OpenStudy (anonymous):

Which equation would best help solve the following problem? Martina hits a volleyball ball upward with an initial velocity of 13 meters/sec. When she hits the ball, it is one meter above the ground. How long will it take the ball to hit the ground?

OpenStudy (anonymous):

A. –4.9t^2 + 13t + 1 = 0 B. –4.9t^2 + 13t = 1 C. 4.9t^2 + 13t + 1 = 0 D. –16t^2+ 13t + 1 = 0